An IQ test was administered to 5 persons before and after training. The results are given below:
             IQ before training: 110      120      123      132          125
             IO after training:   120       118      125       136         121
Test whether there is any change in IQ after training programme. ( t005(4)=2.78)

Grade:12th pass

1 Answers

Arun
6 years ago
-nd)/n-1)=sqrt(140-5(4))/5-1=5.673Hence there is no change of IQ after training2=100+4+4+16+16=140d=sum of d/n=10/2=5 and s=sqrt((d2The results are given below : I II III IV V I.Q before training 110 120 123 132 125I.Q after training 120 118 125 136 121deviation in each case is 10 -2 2 4 -4sum of d=10d